OpenStudy (anonymous):

Simplify completely quantity x squared minus 3 x minus 54 over quantity x squared minus 18 x plus 81 times quantity x squared plus 12 x plus 36 over quantity x plus 6 quantity x plus 6 over quantity x minus 9 the quantity x plus 6 end quantity over the quantity x minus 9 end quantity squared. the quantity x plus 6 end quantity squared over the quantity x minus 9 the quantity x plus 6 end quantity squared over the quantity x minus 9 end quantity squared.

OpenStudy (pfenn1):

Is this the correct expression?\[\frac{x^2-3x-54}{x^2-18x+81}\times \frac{x^2+12x+36}{x+6}\]